OpenStudy (anonymous):

The height (in feet) of a rocket from ground level is given by the function f(t) = -16t2 + 160t. What is the instantaneous velocity of the rocket 3 seconds after it is launched? 32 feet/second 64 feet/second 12 feet/second 10 feet/second 8 feet/second

OpenStudy (michele_laino):

hint: we have to compute the first derivative of s(t), namely: \[\Large v\left( t \right) = \frac{{d\left( { - 16{t^2} + 160t} \right)}}{{dt}} = ...?\]

OpenStudy (michele_laino):

oops..f(t), not s(t)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d is?

OpenStudy (michele_laino):

it is a symbol. I can rewrite that formula, as below: \[\Large v\left( t \right) = \left( { - 16{t^2} + 160t} \right)' = ...?\]

OpenStudy (anonymous):

a 336

OpenStudy (michele_laino):

no, it is: \[\Large v\left( t \right) = \left( { - 16{t^2} + 160t} \right)' = - 32t + 160\] since the first derivative of t^2 is 2t, wheras the first derivative of t, is 1

OpenStudy (michele_laino):

whereas*

OpenStudy (anonymous):

I got a 64

OpenStudy (michele_laino):

here are more steps: \[\Large \begin{gathered} v\left( t \right) = \left( { - 16{t^2} + 160t} \right)' = - 16 \times 2t + 160 \times 1 = \hfill \\ = - 32t + 160 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

the requested velocity, is: \[\Large v\left( 3 \right) = - 32 \times 3 + 160 = 64\;feet/\sec \]

OpenStudy (michele_laino):

so, your answer is right!

OpenStudy (anonymous):

preciate it

OpenStudy (michele_laino):

thanks!